A surface ocean wave has an amplitude of 0.60 m and the distance from trough to trough is 8.00 m. It moves at a constant wave speed of 1.50 m/s propagating in the positive x-direction. At t=0, the water displacement at x = 0 is zero, and vy is negative. Assuming the wave can be modeled as a sine wave, write a wave function to model the wave in 3 ways.

Pls do not forget the unit


1
Expert's answer
2022-05-18T08:32:17-0400
"y(x,t)=-A\sin\left(\frac{2\pi}{\lambda}(x-vt)\right)"


"y(x,t)=-(0.60\:{\rm m})\sin\left(\frac{2\pi}{8.0\:\rm m}(x-1.50t)\right)"